OpenStudy (anonymous):

how do you solve (18-n)^1/2=(n/8)^1/2

OpenStudy (anonymous):

=> sqrt(18-n)=sqrt(n/8) => 18-n=n/8 =>18= n/8+n => 9n/8 =18 => n = 18*8/9

OpenStudy (anonymous):

because to the power 1/2 is actually square root

OpenStudy (anonymous):

what do you have to do with both the 1/2

OpenStudy (anonymous):

x^(1/2)=sqrt(x) understand now ?

OpenStudy (anonymous):

a little =) im looking over the answer just to see how you did it